OpenStudy (anonymous):

I really need help with this. Don't give me the answers though, I want to know how to solve the question, and I want someone to guide me to the answers. Thanks! At times, it is necessary to make plans even when all of the facts are not there. Bernard is currently facing such a dilemma. Bernard is asked to create a swimming pool and deck design for a new series of model homes currently being designed. Each home will have varying pool sizes depending on the lot size; however, there will be some similarities, so polynomial expressions will be used. Your task will be to help Bernard.

OpenStudy (anonymous):

To fit the dimensions of the various land spaces, he realizes that the length of the pool has to be three times the size of the width plus 3, and the depth has to be 7 less than twice the length. Task 1 Create expressions to represent the dimensions of the pool. Create a variable to represent the width of the pool. Determine the expressions to represent length and the depth in terms of the width variable. Determine the polynomial expression to represent the area of land space that the pool will cover. Using the area and the depth expressions, determine the polynomial expression that is used to represent the volume. Classify the volume polynomial by degree and number of terms.

OpenStudy (anonymous):

@triciaal @sleepyjess @Nnesha @ganeshie8 @mathmale @mathstudent55 @zepdrix

OpenStudy (triciaal):

will not have time to complete everything but let's start with a sketch of the pool the length expressed in terms of the width, w the width, w and the depth in terms of the width the pool is like a cuboid, so 3 dimentions volume of pool = length * width * depth

OpenStudy (anonymous):

This is my answer for part A

OpenStudy (anonymous):

a. A variable to represent the width of the pool is x. The length of the pool has to be three times the size of the width + 3. Therefore, the length of the pool would be 3x + 3. The depth has to be 7 less than twice the length. (3x + 3) * 2 = 2(3x + 3) = 6x + 6 - 7 = 6x - 1. Therefore, the width of the pool is x, the length of the pool 3x + 3, and the depth of the pool is 6x – 1.

OpenStudy (anonymous):

Is this correct @triciaal ?

OpenStudy (triciaal):